# MCGA: A Multi-task Classical Chinese Literary Genre Audio Corpus

### v1

@@ Abstract
With the rapid advancement of Multimodal Large Language Models (MLLMs), their potential has garnered significant attention in Chinese Classical Studies (CCS). While existing research has primarily focused on text and visual modalities, the audio corpus within this domain remains largely underexplored. To bridge this gap, we propose the Multi-task Classical Chinese Literary Genre Audio Corpus (MCGA). It encompasses a diverse range of literary genres across six tasks: Automatic Speech Recognition (ASR), Speech-to-Text Translation (S2TT), Speech Emotion Captioning (SEC), Spoken Question Answering (SQA), Speech Understanding (SU), and Speech Reasoning (SR). Through the evaluation of ten MLLMs, our experimental results demonstrate that current models still face substantial challenges when processed on the MCGA test set. Furthermore, we introduce an evaluation metric for SEC and a metric to measure the consistency between the speech and text capabilities of MLLMs. We release MCGA and our code to the public to facilitate the development of MLLMs with more robust multidimensional audio capabilities in CCS. [^1] MCGA Corpus: https://github.com/yxduir/MCGA
@@
Yexing Du1,2   Kaiyuan Liu1,2   Bihe Zhang3   Youcheng Pan2   Bo Yang2   Liangyu Huo4 Xiyuan Zhang4   Xie Jian4   Daojing He1   Yang Xiang2   Ming Liu1,2   Bing Qin1,2 1Harbin Institute of Technology  2Pengcheng Laboratory 3South China University of Technology  4Du xiaoman {yxdu, mliu}@ir.hit.edu.cn, {panych, xiangy}@pcl.ac.cn
@@ 1 Introduction
  • MCGA Corpus: We present MCGA, the first large-scale (119 hours), open-source, and fully copyrighted audio corpus dedicated to Classical Chinese literature. This resource effectively bridges the gap in high-quality audio datasets for this domain.
@@
  • Evaluation Metrics: We introduce 2 novel evaluation metrics: a domain-specific metric tailored for literary SEC, and a Cross-Modal Consistency (CMC) metric designed to assess the alignment between auditory and textual representations.
@@ 3.2 Data Construction
Subsequently, we leverage DeepSeek-V3.2 Guo et al. (2025) to generate question-answer pairs for the text of each clip, with access to the full literary context. This process covers a variety of speech-related tasks, including S2TT, SEC, SU, and SR.
@@
Figure 3: MCGA Corpus Construction. Initially comprising only metadata such as titles, authors, and texts, the MCGA corpus is expanded through human recording, LLM generation, and rigorous verification. Then, it supports six speech tasks: ASR, S2T, SEC, SQA, SU, and SR. We provide a detailed example of the SEC task in Figure 4.
@@ 3.6 Dataset Statistics
The corpus comprises 22,000 filtered human-recorded speech samples (totaling 119 hours) and supports 6 downstream tasks: ASR, S2TT, SEC, SQA, SU, and SR. The longest audio sample is 30 seconds, the shortest is 3.5 seconds, and the average duration is 19.5 seconds.
@@ 4.1 Experiment Setting
GPT-4o-mini-Audio OpenAI (2023) 20.5 42.2 6.0 31.2 75.0 70.4
Gemini-3-Flash Team and DeepMind (2025) 7.0 74.5 54.0 48.7 86.6 83.7
@@
Phi-4-Multimodal-Instruct Abouelenin et al. (2025) 58.2 26.2 12.7 24.5 51.2 54.5
Voxtral-Mini Liu et al. (2025a) 28.0 24.7 14.8 12.4 59.6 62.5
Voxtral-Small Liu et al. (2025a) 29.3 33.2 14.6 28.3 72.9 71.8
MiDashengLM Dinkel et al. (2025) 11.7 43.9 22.6 22.5 72.2 75.6
Step-Audio-2-mini Wu et al. (2025) 10.1 43.0 39.9 45.3 80.7 79.9
Qwen2-Audio-7B-Instruct Chu et al. (2024) 18.9 30.6 24.1 25.1 71.9 64.7
Qwen2.5-Omni-7B Xu et al. (2025a) 10.5 50.5 36.8 43.2 81.3 79.5
Qwen3-Omni-30B-A3B-Instruct Xu et al. (2025b) 4.4 70.0 59.5 51.8 87.1 83.2
@@ 4.2 Main Results
In Table 3, Qwen3-Omni demonstrates superior performance on the MCGA test set for Chinese understanding and generation tasks, specifically in ASR (4.4 CER $\downarrow$), SEC (59.5 LLM-C $\uparrow$), SQA (51.8 F1 $\uparrow$), and SU (87.1 Acc $\uparrow$). Conversely, closed-source models such as Gemini-3-Flash maintain a competitive edge in English generation and Chinese reasoning tasks, leading in metrics such as S2TT (74.5 LLM-B $\uparrow$) and SR (83.7 Acc $\uparrow$). Overall, the open-source models have achieved a competitive level of performance compared with the closed-source ones.
@@ 4.3.1 Analysis of ASR Task
Phi-4-Multimodal-Instruct 58.5 61.3 62.5 60.0 50.7
Voxtral-Mini 27.4 24.7 27.8 32.4 27.1
Voxtral-Small 30.5 24.7 29.6 32.7 28.7
@@
Step-Audio-2-mini 9.0 6.8 7.5 15.1 10.8
Qwen2-Audio-7B-Instruct 19.1 16.7 15.9 23.1 18.7
Qwen3-Omni-30B-A3B-Instruct 3.8 2.9 3.8 6.4 4.6
Qwen2.5-Omni-7B 11.6 7.8 8.8 15.1 8.6
Qwen-Omni-MCGA 2.8 2.9 7.7 5.2 4.2
@@
Figure 7: CER Across Dynasties and Genres. CER refers to $(1-\text{CER}\%)$
ASR Quality.
Table 5 reveals a 0.2 CER gap (Qwen3-Omni) between human-verified valid/test sets and the train set, confirming high data consistency. Residual errors primarily stem from uncommon characters and phonetic loanwords (tongjiazi) in Classical Chinese. Figure 7 shows the CER distribution across dynasties and genres.
@@
Step-Audio-2-mini 10.5 10.1 10.1 10.5
Qwen2.5-Omni-7B 9.8 9.5 10.5 9.8
Qwen3-Omni-30B-A3B-Instruct 4.6 4.4 4.4 4.6
@@ 4.3.2 Analysis of S2TT Task
Phi-4-Multimodal-Instruct 26.0 26.2 27.2 26.5
Voxtral-Mini 24.7 24.7 23.9 24.4
Voxtral-Small 33.3 33.2 32.7 33.1
MiDashengLM 43.0 43.9 40.1 42.4
Step-Audio-2-mini 42.8 43.0 39.8 41.9
Qwen2-Audio-7B-Instruct 29.5 30.6 30.7 30.3
Qwen2.5-Omni-7B 50.0 50.5 45.3 48.6
Qwen3-Omni-30B-A3B-Instruct 68.7 70.0 66.5 68.4
Table 6: Beauty Evaluation of Translation. Following Chen et al. (2025), we employ Beauty of Form (BF), Beauty of Meaning (BM) and Beauty of Sound (BS) as evaluation metrics. LLM-B denotes the mean of three evaluation metrics.
@@ 4.3.3 Analysis of SEC Task
  • Persona Recognition (SEC-P, 0–2): Measures the capability to extract identity features such as age and gender. Starting from an initial score of 2, 1 point is deducted for each attribute error.
@@
Phi-4-Multimodal-Instruct 4.6 7.7 0.4 12.7
Voxtral-Mini 4.9 9.1 0.8 14.8
Voxtral-Small 1.7 10.8 2.1 14.6
MiDashengLM 13.1 6.6 2.9 22.6
Step-Audio-2-mini 16.3 12.6 11.1 39.9
Qwen2-Audio-7B-Instruct 10.2 11.4 2.5 24.1
Qwen2.5-Omni-7B 14.1 13.9 8.8 36.8
Qwen3-Omni-30B-A3B-Instruct 16.1 19.0 24.4 59.5
@@ 4.3.4 Analysis of SQA, SU, and SR Tasks
Phi-4-Multimodal-Instruct 24.5 51.2 54.5 25.1 69.5 61.0 86.9
Voxtral-Mini 12.4 59.6 62.5 13.6 77.5 69.5 86.0
Voxtral-Small 28.3 72.9 71.8 39.6 89.9 83.6 79.5
MiDashengLM 22.5 72.2 75.6 41.7 89.6 85.0 74.5
Step-Audio-2-mini 45.3 80.7 79.9 51.3 89.8 85.8 90.4
Qwen2-Audio-7B-Instruct 25.1 71.9 64.7 35.6 79.6 71.1 83.9
Qwen2.5-Omni-7B 43.2 81.3 79.5 53.6 91.3 84.9 87.8
Qwen3-Omni-30B-A3B-Instruct 51.8 87.1 83.2 60.4 93.6 91.0 90.1
@@ 5 Conclusion
This paper introduces MCGA, the first large-scale, fully copyrighted audio corpus for classical Chinese literature, featuring six speech-language tasks. We develop an evaluation metric for literary SEC and a metric to assess cross-modal consistency. Our systematic evaluation of 10 MLLMs shows that the Qwen-series models demonstrated superior proficiency in understanding CCS.

### v2

@@ Abstract
With the rapid advancement of Multimodal Large Language Models (MLLMs), their potential has gained significant attention in Chinese Classical Studies (CCS). While existing research primarily focuses on text and visual modalities, the audio corpus within this domain remains largely underexplored. To bridge this gap, we introduce the Multi-task Classical Chinese Literary Genre Audio Corpus (MCGA), a 119-hour corpus comprising 22,000 audio samples. It encompasses a diverse range of literary genres across six tasks: Automatic Speech Recognition (ASR), Speech-to-Text Translation (S2TT), Speech Emotion Captioning (SEC), Spoken Question Answering (SQA), Speech Understanding (SU), and Speech Reasoning (SR). Through the evaluation of ten MLLMs, our experimental results demonstrate that current MLLMs still face substantial challenges on the MCGA test set. Furthermore, we introduce a domain-specific metric for SEC and a metric to measure the consistency between speech and text capabilities. We release MCGA to the public to facilitate the development of more robust MLLMs. [^1] MCGA Corpus: https://github.com/yxduir/MCGA
@@
Yexing Du1,2   Kaiyuan Liu1,2   Bihe Zhang3   Youcheng Pan2   Bo Yang2   Liangyu Huo4 Xiyuan Zhang4   Jian Xie4   Daojing He1   Yang Xiang2   Ming Liu1,2   Bing Qin1,2 1Harbin Institute of Technology  2Pengcheng Laboratory 3South China University of Technology  4Du xiaoman {yxdu, mliu}@ir.hit.edu.cn, {panych, xiangy}@pcl.ac.cn
@@ 1 Introduction
  • MCGA Corpus: We present MCGA, the first large-scale (119 hours), open-source, and fully copyrighted audio corpus dedicated to classical Chinese literature studies. This resource effectively bridges the gap in high-quality audio datasets for this domain.
@@
  • Evaluation Metrics: We introduce two novel evaluation metrics: a domain-specific metric tailored for the literary SEC task, and a CMC metric designed to assess the alignment and quantify the gap between auditory and textual modalities.
@@ 3.2 Data Construction
Subsequently, we leverage DeepSeek-V3.2 Guo et al. (2025) to generate question-answer pairs for the text of each clip, with access to the full literary context. This process covers a variety of speech-related tasks, including S2TT, SEC, SQA, SU, and SR.
@@
Figure 3: MCGA Corpus Construction. Initially comprising only metadata such as titles, authors, and texts, the MCGA corpus is expanded through human recording, LLM generation, and rigorous verification. Then, it supports six speech tasks: ASR, S2TT, SEC, SQA, SU, and SR. We provide a detailed example of the SEC task in Figure 4.
@@ 3.6 Dataset Statistics
The corpus comprises 22,000 filtered human-recorded audio samples (totaling 119 hours) and supports 6 downstream tasks: ASR, S2TT, SEC, SQA, SU, and SR. The longest audio sample is 30 seconds, the shortest is 3.5 seconds, and the average duration is 19.5 seconds.
@@ 4.1 Experiment Setting
GPT-4o-mini-Audio OpenAI (2023) 20.6 43.5 5.7 30.6 74.8 70.2
Gemini-3-Flash Team and DeepMind (2025) 6.1 74.0 54.0 48.7 86.6 83.7
@@
Phi-4-Multimodal-Instruct Abouelenin et al. (2025) 59.6 27.5 12.7 24.5 50.6 54.4
Voxtral-Mini Liu et al. (2025a) 30.0 25.5 15.0 12.5 58.9 62.8
Voxtral-Small Liu et al. (2025a) 30.0 34.1 16.4 27.9 72.6 71.9
MiDashengLM Dinkel et al. (2025) 11.7 42.9 24.7 22.5 72.2 75.6
Step-Audio-2-mini Wu et al. (2025) 9.9 41.9 36.8 45.2 80.5 80.4
Qwen2-Audio-7B-Instruct Chu et al. (2024) 18.7 30.4 26.1 24.8 72.1 64.7
Qwen2.5-Omni-7B Xu et al. (2025a) 10.1 49.7 37.0 43.5 81.3 79.3
Qwen3-Omni-30B-A3B-Instruct Xu et al. (2025b) 4.4 67.6 58.4 51.5 86.9 82.9
@@ 4.2 Main Results
In Table 3, Qwen3-Omni demonstrates superior performance on the MCGA test set for Chinese understanding and generation tasks, specifically in ASR (4.4 CER $\downarrow$), SEC (58.4 LLM-C $\uparrow$), SQA (51.5 F1 $\uparrow$), and SU (86.9 Acc $\uparrow$). Conversely, closed-source models such as Gemini-3-Flash maintain a competitive edge in English generation and Chinese reasoning tasks, leading in metrics such as S2TT (74.0 LLM-B $\uparrow$) and SR (83.7 Acc $\uparrow$). Overall, the open-source models have achieved a competitive level of performance compared with the closed-source ones.
@@ 4.3.1 Analysis of ASR Task
Phi-4-Multimodal-Instruct 61.0 63.8 63.0 61.1 51.1
Voxtral-Mini 29.2 27.4 29.2 34.5 28.9
Voxtral-Small 30.6 26.4 30.1 33.9 28.4
@@
Step-Audio-2-mini 9.0 7.0 7.5 14.6 10.3
Qwen2-Audio-7B-Instruct 18.2 15.8 15.9 23.3 19.0
Qwen3-Omni-30B-A3B-Instruct 3.8 2.8 4.1 6.2 4.3
Qwen2.5-Omni-7B 9.9 7.5 8.9 14.8 8.8
Qwen-Omni-MCGA 2.8 3.1 7.8 5.3 4.1
@@
Figure 7: CER Across Dynasties and Genres.
Audio Quality.
Table 5 reveals a 0.1 CER gap (Qwen3-Omni) between human-verified valid/test sets and the train set, confirming high data consistency. Residual errors primarily stem from uncommon characters and phonetic loanwords (tongjiazi) in Classical Chinese. Figure 7 shows the CER distribution across dynasties and genres.
@@
Step-Audio-2-mini 10.4 10.1 9.9 10.4
Qwen2.5-Omni-7B 9.8 9.5 10.1 9.8
Qwen3-Omni-30B-A3B-Instruct 4.5 4.4 4.4 4.5
@@ 4.3.2 Analysis of S2TT Task
Phi-4-Multimodal-Instruct 27.0 27.3 28.1 27.5
Voxtral-Mini 25.6 25.6 25.1 25.5
Voxtral-Small 34.5 34.2 33.6 34.1
MiDashengLM 43.8 44.3 40.6 42.9
Step-Audio-2-mini 42.6 43.3 39.6 41.9
Qwen2-Audio-7B-Instruct 29.8 30.6 30.9 30.4
Qwen2.5-Omni-7B 51.3 51.7 46.1 49.7
Qwen3-Omni-30B-A3B-Instruct 68.1 69.1 65.5 67.6
Table 6: Beauty Evaluation of Translation. Following Chen et al. (2025), we employ Beauty of Form (BF), Beauty of Meaning (BM) and Beauty of Sound (BS) as evaluation metrics. LLM-B denotes the mean of these three evaluation metrics.
@@ 4.3.3 Analysis of SEC Task
  • Persona Recognition (SEC-P, 0–2): Measures the capability to extract identity features such as age and gender. Starting from an initial score of $2$, a penalty of $1$ point is deducted for each attribute error.
@@
Phi-4-Multimodal-Instruct 4.3 7.9 0.5 12.7
Voxtral-Mini 5.5 8.9 0.6 15.0
Voxtral-Small 2.2 11.9 2.3 16.4
MiDashengLM 13.6 8.0 3.2 24.7
Step-Audio-2-mini 16.2 11.8 8.8 36.8
Qwen2-Audio-7B-Instruct 12.9 10.4 2.7 26.1
Qwen2.5-Omni-7B 14.3 13.9 8.8 37.0
Qwen3-Omni-30B-A3B-Instruct 16.0 18.6 23.8 58.4
@@ 4.3.4 Analysis of SQA, SU, and SR Tasks
Phi-4-Multimodal-Instruct 24.5 50.6 54.4 25.2 69.1 60.4 86.8
Voxtral-Mini 12.5 58.9 62.8 13.7 76.8 69.5 86.1
Voxtral-Small 27.9 72.6 71.9 40.0 88.8 83.1 79.3
MiDashengLM 22.5 72.2 75.6 45.1 89.1 84.4 73.5
Step-Audio-2-mini 45.2 80.5 80.4 51.3 89.6 85.4 90.7
Qwen2-Audio-7B-Instruct 24.8 72.1 64.7 36.0 79.9 70.7 83.5
Qwen2.5-Omni-7B 43.5 81.3 79.3 53.1 91.6 85.0 88.0
Qwen3-Omni-30B-A3B-Instruct 51.5 86.9 82.9 60.0 93.6 91.0 89.9
@@ 5 Conclusion
This paper presents MCGA, the first large-scale, fully copyrighted audio corpus dedicated to CCS, encompassing six tasks. Besides, we propose a domain-specific evaluation metric for SEC alongside a metric for cross-modal consistency. Our systematic evaluation of ten MLLMs reveals that the Qwen series leads in performance, exhibiting superior proficiency in comprehending classical Chinese literature.